Answer the Following Problem




Direction: Find all the currents and voltages across each resistor and cell in the following circuits;




1.



































2.









































3.









































4.






































5.









































6.






1
Expert's answer
2021-11-23T10:32:56-0500


for left loop:

"4I_3+3I_4=12"

for right loop:

"2I_2-4I_3=-5"


"I_1=I_2+I_3"

"I_4=I_2+I_3"


then:

"4I_3+3(I_2+I_3)=12"

"I_2=4-7I_3\/3"


"2(4-7I_3\/3)-4I_3=-5"

"I_3=13\cdot3/26=39/26=1.5" A

"I_2=4-7\cdot 1.5/3=0.5" A

"I_1=I_4=0.5+1.5=2" A


"U_2=2\cdot0.5=1" V

"U_3=4\cdot1.5=6" V

"U_4=3\cdot2=6" V